the current ratio may be a misleading meaure of liquidity. true/false and explain
Answer #1

The current ratio may be a misleading measure of liquidity. The given statement is true.

A high current ratio is not always a good indicator of liquidity. High current ratio may be possible due to the presence of less liquid assets such as accrued incomes, prepaid expenses , raw material or work in process inventory. On the other hand, a low current ratio means low liquidity level. But it may also be misleading if current assets mostly consist of cash balance. It can be explained as under:

A Ltd's current assets and current liabilities consist of the following:

Cash $200, Accrued incomes $1,500, Prepaid expenses $2,100, Accounts payable $1,100, Expenses payable $800

Current assets = Cash + Accrued incomes + Prepaid expenses

= 200 + 1,500 + 2,100

= $3,800

Current liabilities = Accounts payable + Expenses payable

= 1,100 + 800

= $1,900

Current ratio = Current assets/Current liabilities

= 3,800/1,900

= 2

B Ltd's current assets and current liabilities consist of the following:

Cash $2,200, Accrued incomes $100, Prepaid expenses $100, Accounts payable $1,100, Expenses payable $500

Current assets = Cash + Accrued incomes + Prepaid expenses

= 2,200 + 100 + 100

= $2,400

Current liabilities = Accounts payable + Expenses payable

= 1,100 + 500

= $1,600

Current ratio = Current assets/Current liabilities

= 2,400/1,600

= 1.5

Current ratio of A Ltd. is 2 while Current ratio of B Ltd. is 1.5. Without examining the components of current ratio, one would say current ratio of A Ltd. is better than current ratio of B Ltd i.e. one would believe that liquidity position of A Ltd. is better than liquidity position of B Ltd.

But an analysis of current assets clearly shows that B Ltd. is more liquid than A Ltd. A Ltd's current assets mostly consist of less liquid assets, it has very less cash available to pay its current liabilities. B Ltd.'s current assets mostly consist of cash, hence its more liquid although its current ratio is less than current ratio of A Ltd.

Hence, current ratio sometimes may be misleading.

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.